### v2

# Optimal Control with Natural Images: Efficient Reinforcement Learning using Overcomplete Sparse Codes

Abstract
Optimal control and sequential decision making are widely used in many complex tasks. Optimal control over a sequence of natural images is a first step towards understanding the role of vision in control. Here, we formalize this problem as a reinforcement learning task, and derive general conditions under which an image includes enough information to implement an optimal policy. Reinforcement learning is shown to provide a computationally efficient method for finding optimal policies when natural images are encoded into “efficient” image representations. This is demonstrated by introducing a new reinforcement learning benchmark that easily scales to large numbers of states and long horizons. In particular, by representing each image as an overcomplete sparse code, we are able to efficiently solve an optimal control task that is orders of magnitude larger than those tasks solvable using complete codes. Theoretical justification for this behaviour is provided. This work also demonstrates that deep learning is not necessary for efficient optimal control with natural images.
Keywords Natural image sequences $\cdot$ Optimal policies $\cdot$ Sufficient statistics $\cdot$ Sparse coding $\cdot$ Benchmark
1 Introduction
Many interesting and complex tasks can be described using natural image sequences. Each image of the sequence corresponds to a particular state of the environment, and choosing an action or applying a control at the current state takes the environment to its next state, and leads to the next image in the sequence. A natural question is then: given an image, what is the best control to apply? Although the answer will depend on the particular task at hand a general framework can be identified. As a simple example, consider moving through a rainforest and taking images of flowers in order to locate a particular species of plant. In this case, each image describes a state of the environment, and the state changes depending on the location of the next image. Each image also has a cost or reward associated with it: an image of the sought-after flower will receive a high reward, while an image without any flowers may receive a low (or even negative) reward. To perform this task well, the choice of control leading to the location of the next image does not necessarily give the best chance of obtaining an image of the sought-after flower at the next time period, rather, it should also lead to a good chance of obtaining such images at future time periods as you move through the rainforest. Planning over a sequence of time periods is the defining characteristic of sequential decision making and optimal control, and is a clear requirement of any intelligent system.
Natural images form a high-dimensional dataset with a great deal of expressive power. For some tasks, images may include enough task-dependent information to determine an optimal policy. In this case no additional information is required for optimal control. In other cases, this might not be true and additional information may be necessary. Here, we will derive the general set of conditions for an image to be sufficient for implementing an optimal policy. Similar to other types of ecological data, it is well-known that the statistical properties of natural images distinguish them from artificially generated images (Field1; Field2; ruderman1; ruderman2; simoncellio; hyvarinenbook; Hosseini). We exploit these statistical properties to find efficient image representations called sparse codes. Sparse codes will then be shown to significantly reduce the computational resources required for finding optimal policies.
The aim of this work is to investigate the possibility of efficient optimal control over a sequence of natural images. The investigation has three stages. The first stage identifies what information an image must include so that optimal control is possible when an image of the environment is the only information available; in other words, when does an image become a “sufficient statistic”? The second stage identifies how to achieve efficient optimal control. Statistical dependencies in natural images lead to redundancy, making it possible to obtain an efficient representation using an appropriate encoding scheme such as sparse coding (daugman3; daugman2; olshausenfield1; olshausenfield2). A sparse code is “efficient” in the sense of being a low bit-rate (low entropy) representation of a high-dimensional dataset. Here, we show an overcomplete sparse code is also computationally efficient for finding optimal policies in reinforcement learning. The basic idea is described by the neural network shown in Figure 1. The first two layers of the neural network form a sparse autoencoder that generates an overcomplete sparse code by reconstructing an image input using an overcomplete basis of Gabor functions adapted to natural image statistics (loxley1). It is well-known that Gabor functions can be used to provide a “sparse description” of natural images (daugman3; daugman2). A linear regression network then makes use of this overcomplete sparse code to approximate the cost-to-go required in reinforcement learning. The efficiency gain is twofold: an overcomplete sparse code is shown to increase both the learning rate and the storage capacity of a linear network used for approximating cost-to-go values. The underlying theoretical reason is that a sparse code favourably conditions the Hessian matrix of the least squares problem, while an overcomplete sparse code also increases the rank of the associated design matrix. The third stage of the investigation empirically proves that an overcomplete sparse code gives the most efficient image representation out of several likely alternatives. This requires developing a new reinforcement learning benchmark to compare the performance and efficiency of different image representations. The benchmark must be scalable so the number of states and time periods can easily be increased; and optimal solutions must easily be distinguishable from suboptimal solutions. Scalability of the benchmark is necessary for determining which image representations will perform best over long horizons and for large numbers of states.
[…]
Figure 1: Neural network for reinforcement learning with natural images. The first two layers form a sparse autoencoder that generates an overcomplete sparse code $\phi$ by reconstructing the image input $I$ using an overcomplete basis of Gabor functions adapted to natural image statistics. The output $r\boldsymbol{\cdot}\phi$ then approximates the cost-to-go $\beta$ using weights $r$. The network storage capacity has increased from $d$ to (close to) $m$ by using an overcomplete sparse code.
Previous work on reinforcement learning is heavily concentrated on synthetic datasets of limited complexity (although the tasks themselves may have very high complexity). Researchers have previously suggested that more work should be done to extend the paradigm of reinforcement learning to real-world data such as natural images and natural video (zhang2018natural). The few available examples of these efforts include using image representations to integrate unsupervised training of deep neural networks into reinforcement learning; specifically for visual navigation (lange), and simulated autonomous driving tasks (zhang2021learning). A little more work exists on the role of sparse codes and sparse representations in reinforcement learning, such as work to obtain better policies with neural networks on certain standard classic control tasks (Martha1; Martha2; Rafati). This line of work dates back to early research on coarse coding in reinforcement learning. One the other hand, none of these works considered sparse codes of images or other high-dimensional datasets of high complexity. The closest previous work appears in loxley2, where sparse representations of images were applied to find optimal policies for a specific control task. Unfortunately, this work was limited by the size of the task; which required a hand-annotated video dataset, and the type of task; which was restricted to deterministic controllable dynamics rather than a general Markov decision process. These limitations did not allow for results to be extended to large numbers of states and long horizons. In this work, a more general approach to optimal control is taken, and the development of a new reinforcement learning benchmark now allows results to be easily scaled up.
The structure of this paper is as follows. In Sec. 2, reinforcement learning methodology is briefly summarized, and the problem of optimal control over a sequence of natural images is formalized. Following this, a new reinforcement learning benchmark is developed in Sec. 3. In Sec. 4, results for the reinforcement learning benchmark are presented, and alternative image representations are compared. A detailed discussion of results and conclusions is given in Sec. 5.
2 Reinforcement Learning over Natural Image Sequences
We consider a general optimal control task involving a discrete time dynamical system with a finite number of states and controls. During any time period $k$, the environment is described by a state $i$ taken from a state space $S$, and a control $u$ is selected from a set of available controls $U_{k}(i)\subset C$ (we adopt the notation from bert). The state of the environment changes when a controller or agent chooses control $u$ from state $i$, moving the environment to state $j$ according to the transition probabilities $p_{ij}(u)$ of a controllable Markov chain. Since the controller or agent can choose different controls depending on the current state of the environment, the controller’s choice is described by a policy $\pi=\{\mu_{0},...,\mu_{N-1}\}$, where the sequence of functions $\mu_{k}$ map states into controls: $\mu_{k}(i)=u$, for each state $i$ and each time period $k$. The cost of being in state $i_{k}$ during time period $k$, choosing control $u_{k}$, then moving to state $i_{k+1}$ during time period $k+1$, is denoted by $g_{k}(i_{k},u_{k},i_{k+1})$. The key quantity of interest is the expected total cost. Given a policy $\pi$ and an initial state $i_{0}$, the expected total cost is the expected value of the sum of costs:
$J_{\pi}(i_{0})={\mathbb{E}}\left\{g_{N}(i_{N})+\sum_{k=0}^{N-1}g_{k}(i_{k},\mu_{k}(i_{k}),i_{k+1})\right\},$ (1)
where $g_{N}(i_{N})$ is the terminal cost incurred at the end of the task ($k=N$); and the expectation is taken over the states $i_{1},...,i_{N-1}$ using the transition probabilities $p_{ij}(\mu_{k}(i))$, the initial state $i_{0}$, and the chosen policy $\pi=\{\mu_{0},...,\mu_{N-1}\}$. The goal of optimal control is to choose a policy so as to minimize the expected total cost; i.e.,
$\text{minimize}\ \ \ J_{\pi}(i_{0}),$ (2)
$\text{subject to}\ \ \ \mu_{k}(i)\in U_{k}(i)\ \ \ \ \ \ \forall k,i.$
The finite-horizon optimal control task outlined in Eqs. (1) and (2) can be solved exactly using the Dynamic Programming Algorithm, leading to the optimal cost-to-go $J_{k}^{*}$ (this algorithm is called Value Iteration in the case of infinite horizon tasks, and converges to $J^{*}$ for discounted problems with stationary policies). The optimal policy $\mu_{k}^{*}(i)$ is then given by the minimizing control $u^{*}$ in the expression:
$\underset{u\in U_{k}(i)}{\operatorname{min}}\ \sum_{j}p_{ij}(u)\left(g_{k}(i,u,j)+J_{k+1}^{*}(j)\right),$ (3)
for each state $i$ and each time period $k$.
Optimal control with natural image sequences requires a slight modification of this general framework. The key consideration is now what information to include in a state to allow for optimal decision making. Any state that includes all necessary information for optimal control is called a sufficient statistic (bert). The assumption we make here is that an image is a sufficient statistic. In this case, each state $i\in S$ in an optimal control task is now given by an image $\phi(i)\in\bar{S}$; where $\bar{S}\subset\mathbb{R}^{p}$ so that each image is represented as a $p$-dimensional vector. The index $i$ used to denote a particular state in the general framework now denotes a particular image (state) in the modified framework. However, this index is simply a notational convenience to distinguish between different images, and is not part of the information required for optimal control. An image $\phi(i)$ is a sufficient statistic if $U_{k}(i),p_{ij}(u)$, and $g_{k}(i,u,j)$ in Eq. (3) depend on $i$ or $j$ only through $\phi(i)$ or $\phi(j)$, respectively. For example, a striking consequence of $\phi(i)$ as a sufficient statistic is the form the transition probabilities now take:
$p_{ij}(u)=\bar{p}(\phi(j)|\phi(i),u),$ (4)
so that given image $\phi(i)$, selecting control $u$ leads to image $\phi(j)$ with probability $\bar{p}(\phi(j)|\phi(i),u)$. Therefore, a controllable Markov chain now becomes an “image generator” when used to generate new samples. We derive a simple “image generator” in the next section. The main consequence of $\phi(i)$ as a sufficient statistic is that the optimal policy can now be written as
$\mu_{k}^{*}(i)=\bar{\mu}_{k}^{*}(\phi(i)),$ (5)
and so optimal control depends only on the information contained in the image $\phi(i)$.
Now that the general framework for optimal control has been appropriately modified to include natural image sequences, we present an algorithm for finding an optimal policy. Given an image $\phi(i)$ we require a function $\bar{J}_{k}(\phi(i))$, analogous to $J_{k}^{*}(i)$, that determines its cost-to-go. A simple choice of function is a linear network given by the scalar product of a weight vector $r_{k}\in\mathbb{R}^{p}$ and an image vector $\phi(i)$:
$\bar{J}_{k}(\phi(i),r_{k})=r^{\top}_{k}\phi(i),$ (6)
where $\top$ is the vector transpose. To approximate the optimal cost-to-go $J_{k}^{*}$ requires adjusting the network weights using a training rule and a dataset. Here, we apply the algorithm known as Fitted Value Iteration. To start the iteration, we go to the end of the task at k=N1k=N-1, and make use of the known terminal cost $\bar{J}_{N}(\phi(i),r_{N}^{*})=g_{N}(i)$. The following pair of expressions are then iterated backwards from k=N1k=N-1 to $k=0$ using $\bar{J}_{N}$ to start the iteration,
$\beta_{k}^{i}=\underset{u\in\bar{U}_{k}(\phi(i))}{\operatorname{min}}\ \sum_{j}\bar{p}(\phi(j)|\phi(i),u)\left\{\bar{g}_{k}(\phi(i),u,\phi(j))+\bar{J}_{k+1}(\phi(j),r_{k+1}^{*})\right\},\ \ \ \ \ \forall i$ (7)
$\text{minimize}\ \ \ \sum_{i}(\bar{J}_{k}(\phi(i),r_{k})-\beta_{k}^{i})^{2},$ (8)
$\text{subject to}\ \ \ r_{k}\in\mathbb{R}^{p},$
yielding the cost-to-goes $\bar{J}_{N-1},...,\bar{J}_{0}$ required for finding an optimal policy. The first expression (7) yields the cost-to-go $\beta_{k}^{i}$ of image $\phi(i)$ at time period $k$ by making use of the cost-to-go $\bar{J}_{k+1}(\phi(j),r_{k+1}^{*})$ of image $\phi(j)$ at time period $k+1$, as well as the current cost $\bar{g}_{k}(\phi(i),u,\phi(j))$. In the second expression (8), the “data” given by $(\phi(i),\beta_{k}^{i})$ from Eq. (7) is used to fit the linear network $\bar{J}_{k}(\phi(i),r_{k})$ to the cost-to-go $\beta_{k}^{i}$ by adjusting the weights $r_{k}$ to minimize the sum of squares.
Natural images used for optimal control in (6)–(8) are computationally inefficient when using raw image pixels, as will be demonstrated in the next section. It turns out to be much better to use some type of image representation. The sparse autoencoder in Figure 1 generates one possible representation for natural images, where $\phi(i)\in\mathbb{R}^{p}$ is given by an overcomplete sparse code (i.e., $p=m>d$). In this case, training the linear network for optimal control requires solving two linear least-squares problems: 1) minimize $\lVert G\phi-I\rVert_{2}^{2}$ to find an overcomplete sparse code $\phi\in\mathbb{R}^{m}$ given an image input $I\in\mathbb{R}^{d}$ and an overcomplete basis of Gabor functions $G\in\mathbb{R}^{d\times m}$ (adapted to natural image statistics), followed by 2) minimize $\lVert r\boldsymbol{\cdot}\phi-\beta\rVert_{2}^{2}$ to find a weight vector $r\in\mathbb{R}^{m}$ after making use of (7) to find $\beta$. This must be done for each image input.
3 A Scalable Benchmark for Efficient Image Representations
We would now like to design a simple and robust benchmark for the purpose of assessing alternative image representations for optimal control with natural image sequences. The number of states of the benchmark should scale to arbitrarily large numbers, and its cost structure and dynamics should be easy to implement over long horizons.
The number of states of the benchmark can scale to arbitrarily large numbers by working with image patches. An image patch is an arbitrarily sized image region extracted from a larger parent image. Discretizing a parent image into a regular grid pattern yields a set of image patches $\bar{S}$, with each image patch $\phi(i)\in\bar{S}$ indexed by $i\in\mathbb{Z}^{2}$: a unique pair of coordinates giving the location of the image patch within its parent image (for example, each coordinate pair could give the coordinates of the top-left corner of an image patch). Given a parent image of fixed size the number of image patches can be increased simply by decreasing the size of each image patch. The following proposition will be used to help construct the benchmark.
Proposition 1.
A state of the benchmark can be described by $i$ or $\phi(i)$, and either state is a sufficient statistic.
Proof.
Given any image patch $\phi(i)\in\bar{S}$, it is possible to determine its index $i$ by running through all image patches to find the closest match:
$\hat{i}=\underset{m}{\operatorname{argmin}}\lVert\phi(m)-\phi(i)\rVert_{2}^{2},$
provided each image patch is unique. This index also corresponds to the state $i\in S$. Alternatively, given state $i$, it is possible to determine the image patch $\phi(i)$ by looking up the $i$th element of $\bar{S}=\{\phi(1),...,\phi(n)\}$. The same is true for both $j$ and $\phi(j)$. Therefore, given $p_{ij}(u)$, $g_{k}(i,u,j)$, and $U_{k}(i)$; it is possible to determine $\bar{p}(\phi(j)|\phi(i),u)=p_{ij}(u)$, $\bar{g}_{k}(\phi(i),u,\phi(j))=g_{k}(i,u,j)$, and $\bar{U}_{k}(\phi(i))=U_{k}(i)$. Alternatively, given $\bar{p}(\phi(j)|\phi(i),u)$, $\bar{g}_{k}(\phi(i),u,\phi(j))$, and $\bar{U}_{k}(\phi(i))$; it is possible to determine $p_{ij}(u)=\bar{p}(\phi(j)|\phi(i),u)$, $g_{k}(i,u,j)=\bar{g}_{k}(\phi(i),u,\phi(j))$, and $U_{k}(i)=\bar{U}_{k}(\phi(i))$. ∎
The proof of Proposition 1 describes a very simple image generator $\phi(j)\sim\bar{p}(\boldsymbol{\cdot}|\phi(i),u)$ constructed from an image patch set $\bar{S}$, and any set of transition probabilities $p_{ij}(u)$. A non-trivial model for $p_{ij}(u)$ is presented in the next sections.
The cost structure and transition probabilities of the benchmark are motivated by dragonfly tracking in videos (loxley2). The tracking task requires a tracker to follow a target as closely as possible over many time periods using a set of discrete controls that determines the dynamics of the tracker (see Figure 2). Letting the image patch coordinates be related to the distance between the tracker and target allows this distance to be used to assign a cost to each image patch. Moving from one image patch to another changes this distance, and therefore the cost. An optimal policy corresponds to the optimal choice of controls for tracking the target starting from any image patch during any time period. This task is easily scalable to long horizons and many states provided the target dynamics scales well. Next, a scalable model for the target dynamics is presented.
3.1 Target Dynamics
The target dynamics for the benchmark has two key requirements: 1) the target must be difficult enough to track so that simple algorithms such as greedy search perform suboptimally, and 2) the dynamics must easily scale to long horizons and many states. Both points can be addressed by introducing a simple generative model for the target dynamics.
Motivated by the dragonfly video dataset used for optimal control in loxley2, the following type of scenario is proposed. The target (T) starts on the bottom row of a $3\times 3$ grid (shown in Fig 1, first grid). The target and the tracker can both move up, left, or right at each time period. However, the target (but not the tracker) can also make an “evade” move by moving diagonally to evade the tracker. The tracker must therefore carefully plan its sequence of moves in order to follow the target closely.
During the first time period (second grid in Fig 1) the target remains where it is, while at the beginning of the second time period (third grid in Fig 1) it makes a diagonal move. An optimal tracker (blue square) anticipates a diagonal move immediately following a time period where the target is stationary, and pays a small cost to move up one square away from the target during the first time period (second grid in Fig 1). Whether the target next moves diagonally left or diagonally right, the optimal tracker is now able to reach the target during the next time period (third grid in Fig 1). On the other hand, a tracker applying a greedy algorithm (red squares) tries to stay as close to the target as possible during each time period, allowing the target to “evade” the greedy tracker when it makes its diagonal move, causing the greedy tracker to fall one move behind the target (third and fourth grids in Fig 1). A greedy tracker will always be suboptimal for this sequence.
T
T
T
T
Figure 2: A target tracking sequence (from left to right) showing optimal and suboptimal trackers. A tracker can move either “up” or “right”, while the target can move “up”, “right”, or “diagonally”. The suboptimal tracker (red square) follows the target (T) as closely as possible at each step (a greedy approach), causing it to fall behind when the target moves diagonally. The optimal tracker (blue square) follows the target by anticipating a diagonal move following a time period where the target is stationary.
The sequence of target moves in Fig 1 can be generalized and modelled as a regular language (sipser). The alphabet of target-moves is $\{\mathrm{s,d,r}\}$; where $\mathrm{s}\rightarrow(0,0)$, $\mathrm{d}\rightarrow(1,1)\ \text{or}\ (-1,1)$, and $\mathrm{r}\rightarrow(0,1)\ \text{or}\ (-1,0)\ \text{or}\ (1,0)$. Each symbol in the alphabet represents a change in the target coordinates, $\Delta t$, as the target moves from its current position to its next position. The regular expression given by $\mathrm{r}^{*}\cup(\mathrm{sdr})^{*}$ generates all strings in this language, representing all possible target dynamics. For example, the string given by $\mathrm{sdr}$” represents the sequence in Fig 1: $\Delta t_{1}=(0,0)$; $\Delta t_{2}=(1,1)$; and $\Delta t_{3}=(0,1)$. According to this sequence, a target starting at $(2,1)$ stays at $(2,1)$ during the first time period; then moves to $(3,2),$ and $(3,3)$ in the second, and third time periods, respectively (see Fig 1).
A Markov chain can be constructed to generate stochastic dynamics obeying this language. The Markov chain is parameterized by the single degree of freedom available in the regular expression: whether a symbol of a substring begins with an $\mathrm{r}$” or an $\mathrm{s}$”. This is parameterized using the probability $p$, leading to the Makov chain shown in Fig 2.
rppd11s1p1-p11
Figure 3: A Markov chain for generating the target dynamics shown in Fig 1. In Fig 1, the states $\mathrm{s,d,}$ and $\mathrm{r}$ correspond to “same position”, “diagonal move”, and “up” or “right”, respectively. The non-zero transition probabilities are: $p(\Delta t_{k}=\mathrm{d}|\Delta t_{k-1}=\mathrm{s})=1$, $p(\Delta t_{k}=\mathrm{r}|\Delta t_{k-1}=\mathrm{d})=1$, $p(\Delta t_{k}=\mathrm{s}|\Delta t_{k-1}=\mathrm{r})=1-p$, and $p(\Delta t_{k}=\mathrm{r}|\Delta t_{k-1}=\mathrm{r})=p$.
Different values of $p$ give different steady state distributions of the Markov chain. When $p=0$ or 1, the target dynamics is a deterministic periodic sequence: $(\mathrm{sdr})^{*}$ for $p=0$, and $\mathrm{r}^{*}$ for $p=1$. When $0<p<1$, the Markov chain is recurrent and non-periodic: i.e., when $p=0.5$ the steady state probabilities of $\mathrm{s,d,}$ and $\mathrm{r}$ are $1/4$, $1/4$, and $1/2$, respectively; and a typical sequence of twelve target moves looks like $\mathrm{sdrrsdrrrrsdr}$.
3.2 Target Tracking
It is now possible to specify the reinforcement learning benchmark as a target tracking task. This task requires two dynamical variables $t_{k},c_{k}\in\mathbb{Z}^{2}$ representing the discrete two-dimensional coordinates of a target, and a controller (tracker), respectively, at time period $k$. Changes to the environment are described by the discrete dynamical system:
$t_{k+1}=t_{k}+\Delta t_{k},$ (9)
$c_{k+1}=c_{k}+u_{k}.$ (10)
A target updates its position from $t_{k}$ to $t_{k+1}$ according to the transition probabilities of the Markov chain presented in Sec. 3.1. In order to follow the target as closely as possible a controller updates its position from $c_{k}$ to $c_{k+1}$ by choosing a control $u_{k}$ from the set $U$ of available controls. The state information required for choosing an optimal control is given by defining the state to be:
$i_{k}=(c_{k}-t_{k},\Delta t_{k-1})\in D\times T,$ (11)
where $c_{k}-t_{k}\in D$, and $\Delta t_{k-1}\in T$; so that $S=D\times T$. Here, $D,T,U\subset\mathbb{Z}^{2}$ are sets of integer pairs. In order to derive the transition probabilities $p_{ij}(u)$, we proceed as follows. According to Eq. (11), and making use of Eqs. (9) and (10), the state update is given by
$i_{k+1}$ $=(c_{k+1}-t_{k+1},\Delta t_{k}),$
$=(c_{k}-t_{k}+u_{k}-\Delta t_{k},\Delta t_{k}).$ (12)
Denoting the state $i=i_{k}$ and its successor $j=i_{k+1}$ in terms of their individual components:
$i=(a_{1},b_{1}),\ \ \ \ j=(a_{2},b_{2}),$ (13)
and upon equating with Eqs. (11) and (12), leads to $a_{1}=c_{k}-t_{k},b_{1}=\Delta t_{k-1},a_{2}=a_{1}+u_{k}-\Delta t_{k},$ and $b_{2}=\Delta t_{k}$. These equations can be viewed as constraints on the allowed values of $j$ via $a_{2}$ and $b_{2}$, and represented as $\delta(a_{2}-(a_{1}+u_{k}-\Delta t_{k}))$ and $\delta(b_{2}-\Delta t_{k})$; where $\delta(k)=1$ if $k=0$, and $\delta(k)=0$ otherwise. The controllable Markov chain is governed by the the Markov chain from Fig. 3 describing the target dynamics: $p(\Delta t_{k}|\Delta t_{k-1})$, as well as the constraints on $j$, giving:
$p_{ij}(u)$ $=\sum_{\Delta t_{k}}p(\Delta t_{k}|b_{1})\delta(a_{2}-(a_{1}+u-\Delta t_{k}))\delta(b_{2}-\Delta t_{k}),$ (14)
$=p(b_{2}|b_{1})\delta(a_{2}-(a_{1}+u-b_{2})).$ (15)
The cost per time period is the distance between the tracker and the target at each time period, and is given by
$g_{k}(i,u,j)=\lVert a_{1}\rVert_{2}^{2},$ (16)
where $\lVert.\rVert_{2}$ is the Euclidean distance. Substituting these choices into Expression (3), then making use of Proposition 1 to replace $J_{k+1}^{*}(j)$ with $\bar{J}_{k+1}(\phi(j),r_{k+1}^{*})$, leads to:
$\underset{u\in U}{\operatorname{min}}\ \sum_{b_{2}}p(b_{2}|b_{1})\bar{J}_{k+1}(\phi(a_{1}+u-b_{2},b_{2}),r_{k+1}^{*}).$ (17)
The expressions (7) and (8) are used for determining $r_{k}^{*}$ with (6) used for $\bar{J}_{k}$. The sets $T$ and $U$ are specified by $T=\{(0,0),(1,1),(0,1)\}$ and $U=\{(0,0),(1,0),(0,1)\}$; restricting the target moves as in $T$ leads to a tracking task with less uncertainty and better optimal policies. The set $D$ determines the number of coordinate differences (i.e., distances) between the target and the controller, and depends on the number of states (image patches) chosen for the benchmark. Once $D$ has been specified, an optimal policy for the benchmark is found by carrying out the minimization in (17) at each state $(a_{1},b_{1})\in D\times T$ and at each time period $k$.
4 Results
Results for the reinforcement learning benchmark developed in Section 3 are now presented. All code for the following figures is supplied at Loxley4. We begin by looking at the optimal and greedy policies for target tracking.
4.1 Optimal and Greedy Policies in the Reinforcement-Learning Benchmark
Evasive target dynamics in the benchmark leads to different policies for optimal and greedy tracking. When $p=0$ in the Markov chain, the target dynamics is deterministic and periodic and given by the repeating sequence $(\mathrm{sdr})^{*}$. The optimal policy from Eq. (17) is given in Table 1 when the chain is initially in state $\mathrm{s}$, and cycles between three states: $i_{k}=((0,1),(0,0))$, $i_{k+1}=((0,0),(1,1))$, and $i_{k+2}=((0,0),(0,1))$. Remember, the first pair of each state is the coordinate difference between the target and controller (tracker), while the second pair is the target’s change in coordinates during the previous time period.
State Control Cost
$((0,1),(0,0))$ $(1,0)$ 1
$((0,0),(1,1))$ $(0,1)$ 0
$((0,0),(0,1))$ $(0,1)$ 0
Table 1: Optimal policy $\mu^{*}$ and cost per time period $g$.
Only the first state has a non-zero cost, and therefore the total cost of the optimal policy increases by one every three time periods. This is shown in Fig 4 for $p=0$.
Figure 4: Expected total cost of the optimal policy from Table 1 (blue circles), and the greedy policy from Table 2 (red pluses), as the horizon goes from $N=1$ to 30 time periods for $p=0$ and $p=1$.
The greedy policy can be found by replacing $\bar{J}_{k+1}$ in Eq. (17) with $g_{k+1}$, leading to the simpler expression:
$\underset{u^{G}\in U}{\operatorname{min}}\ \sum_{b_{2}}p(b_{2}|b_{1})\lVert a_{1}+u^{G}-b_{2}\rVert_{2}^{2}.$ (18)
According to Eq. (18), a greedy tracker chooses $u^{G}$ as close as possible to $b_{2}-a_{1}$ (so that $c_{k+1}$ is as close as possible to $t_{k+1}$ at the next time period), without regard for what may happen at future time periods. That is, a greedy tracker has no scope for planning ahead. The greedy policy from Eq. (18) is given in Table 2.
State Control Cost
$((0,0),(0,0))$ $(1,0)$ 0
((0,1),(1,1))((0,-1),(1,1)) $(0,1)$ 1
((0,1),(0,1))((0,-1),(0,1)) $(0,1)$ 1
Table 2: Greedy policy $\mu^{G}$ and cost per time period $g$.
The greedy policy also cycles between three states, with a total cost that increases by two every three time periods, as shown in Fig 4 for $p=0$. This means the greedy policy has twice the total cost of the optimal policy when $p=0$, over any number of complete cycles.
When $p=1$, the target dynamics is given by the repeating sequence $\mathrm{r}^{*}$. The optimal policy in Table 1 is now $\mu^{*}((0,0),(0,1))=(0,1)$, with cost per time period $g((0,0),(0,1))=0$. Similarly, the greedy policy in Table 2 is now $\mu^{G}((0,-1),(0,1))=(0,1)$, with cost per time period g((0,1),(0,1))=1g((0,-1),(0,1))=1. The ratio of total costs of the greedy to optimal policies is therefore N1N-1 to 1, so the greedy policy always has a cost N1N-1 times larger than that of the optimal policy, as seen in Fig. 4 for $p=1$.
When $0<p<1$, the target dynamics is stochastic, recurrent, and non-periodic: at the next time period, state $\mathrm{r}$ has probability $p$ of staying in the same state, and probability 1p1-p of going to state $\mathrm{s}$. The expected total cost of the optimal policy in Table 1 now lies somewhere (not shown) between the curves for the optimal policies with $p=1$ and $p=0$ in Fig. 4. Similarly, the expected total cost of the greedy policy in Table 2 now lies (not shown) between the curves for the greedy policies with $p=0$ and $p=1$ in Fig. 4.
The optimal policy and its expected total cost were found using both fitted value iteration and the dynamic programming algorithm, with both approaches yielding the same result. Results for the greedy policy were found directly from Eq. (1) with $\mu_{k}(i)=\mu_{k}^{G}(i)$ from Eq. (18) using an efficient algorithm described in the Appendix. These results were also confirmed using policy evaluation; that is, by solving $J_{k}(i)=\sum_{j}p_{ij}(\mu_{k}^{G}(i))(g_{k}(i,\mu_{k}^{G}(i),j)+J_{k+1}(j))$ using dynamic programming iterations. In this case, $J_{N}(i)$ was set to the terminal cost as for fitted value iteration and dynamic programming.
4.2 Benchmark Performance with Length of Horizon
The optimal and greedy policies given in Tables 1 and 2 turn out to be stationary policies for all values of $p$. This is confirmed using an infinite horizon analysis. The cost of each policy is found in the Appendix by introducing a discount factor and applying policy evaluation in the limit of an infinite horizon (i.e., $N\rightarrow\infty$). These calculations also agree with value iteration applied to the discounted problem (not shown). Furthermore, the cost ratio of the greedy to optimal policies in Tables 1 and 2 is given in the Appendix in the limit of an infinite horizon for all values of $p$. The result is completely consistent with our discussion of these policies for finite horizons. For example, this ratio becomes 2 when $p=0$, 5 when $p=0.75$, and diverges when $p=1$. For finite horizons, the first result ($p=0$) is most easily seen in Fig 4 when $N=30$; giving a ratio of $20/10=2$. The final result ($p=1$) is also seen in Fig 4 when $N=30$, giving a ratio of $29/1=29$. As $N$ increases, this numerator will increase, diverging in the limit as $N\rightarrow\infty$. The result for $p=0.75$ can be seen from the ratio of expected total costs of the greedy and optimal results in Fig 8.
The optimal and greedy policies lead to more complicated dynamics over shorter horizons when different initial states are chosen. For an initial state given by $((0,1),(0,0))$ (i.e., where the tracker starts close to the target) the bottom pair of curves in Fig. 5 reach the stationary optimal and greedy policies in Tables 1 and 2 at very short horizons. As the initial distance between the tracker and target increases for positive coordinates (middle and top pairs of curves in Fig. 5), the optimal and greedy policies take longer and longer to reach these stationary policies. For example, the top pair of curves has not reached their stationary policies until the horizon is around 14 time periods. These results suggest the dynamics of all initial states with one positive coordinate and one non-negative coordinate will eventually reach the stationary policies given in Tables 1 and 2 provided the horizon is long enough.
Figure 5: Expected total cost of optimal and greedy policies as the horizon goes from $N=1$ to $N=30$ time periods for $p=0$ with different initial states. The bottom pair of curves start at state $((0,1),(0,0))$; the middle pair start at state $((3,0),(0,0))$; and the top pair start at state $((4,0),(0,0))$.
In Fig. 5, the greedy policy is shown to be an optimal policy when the initial state of the tracker is far from the target. Reaching a stationary policy only happens when the tracker gets close enough to the target (i.e., within a distance of one). Once the tracker is close to the target the greedy policy is no longer optimal (as shown at longer horizons in Fig. 5). For initial states where both coordinates are negative, the tracker is always behind the target and can never catch up. A simple greedy approach is then sufficient to follow the target, ensuring the greedy policy is an optimal policy in this case. A more detailed discussion of initial state dependence is given in Sec. 4.5.
4.3 Efficient Image Representations for Reinforcement Learning
Image representations for natural image sequences used in reinforcement learning were found by transforming image patches from the benchmark into overcomplete sparse codes and whitened complete codes. A set of 48 natural images was taken from the database “Natural Scene Statistics in Vision Science” by geisler. Each image was cropped to $2844\times 2844$ pixels, then converted to grayscale and double precision using the Matlab functions rgb2gray and im2double, before being discretized into square image patches of side-length $a$ (measured in pixels). The number of image patches in a single image was therefore $\lfloor 2844/a\rfloor^{2}$; where the function $\lfloor.\rfloor$ rounds down to the nearest integer.
Overcomplete sparse codes were constructed using the sparse autoencoder described in Figure 1 and detailed in the Appendix (see also loxley2). This method relies on the two-dimensional Gabor function, originally used to model simple-cell receptive field profiles in the primary visual cortex (hubel; jones; daugman1; daugman2). The Gabor function parameters were chosen by drawing samples from a tractable multivariate probability distribution that well-approximates natural image statistics (loxley1). The size of the sparse code depends on the number of Gabor functions in the overcomplete basis used to reconstruct an image input. To generate a $\times 64$ overcomplete sparse code, the number of Gabor functions required is $64\times a^{2}$ for an image patch of $a^{2}$ pixels. In order to manage the storage capacity of our network we enforce the inequality $m>n$; where $m$ is the size of an overcomplete sparse code, and $n$ is the number of images (i.e., the number of cost-to-goes stored in the network). For this reason, we choose $a$ to maximize the number of image patches available in a $2844\times 2844$ image, whilst ensuring the size of a $\times 64$ overcomplete sparse code is larger than this number; i.e., $64\times a^{2}>\lfloor 2844/a\rfloor^{2}$. The unique solution to this problem is $a=19$, leading to 22201 image patches of 361 pixels each.
A more traditional method of pixel decorrelation (whitening) was used to construct complete codes: where the representation has the same number of pixels as the image patch it was derived from. This method was carried out by setting the mean pixel values (taken over all image patches) to zero, and diagonalizing the corresponding covariance matrix. This method does not generalize to overcomplete codes.
4.4 Benchmark Performance with Number of States
Training a neural network on efficient image representations takes fewer computational resources than training it directly on images (loxley2). This can be seen in Figs. 6 and 7, where the number of least squares iterations required to store a given number of cost-to-go values in a linear network is shown for network inputs given by image patches, and representations of image patches. Each state of the benchmark is given by a unique image patch that is associated with a cost-to-go value. By adjusting the network weights during training, this association is stored in the network in order to determine an optimal policy. In a linear network, the number of network weights determines the storage capacity of the network; that is, the maximum number of cost-to-go values it can store. In turn, the number of weights is equal to the number of pixel inputs to the network; since each input is multiplied by its corresponding weight and summed together to yield the network output. Therefore, the storage capacity of a network can be increased by increasing the number of pixel inputs and network weights. One way to do this for a fixed-sized image patch is to generate an overcomplete representation of an image patch, and use this as input to the network instead – as shown in Figure 1. This approach will be used to solve the benchmark tracking task as the number of states and stored cost-to-go values increases. Network training is carried out using the Matlab lsqr function running on an Nvidia GPU, with the number of least squares iterations until convergence (within $1\times 10^{-6}$ of the least squares objective), averaged over 48 independent trials, shown along the vertical axis.
Figure 6: Number of least squares iterations to store a number of cost-to-go values in a linear network using different image patch representations (see legend). The least squares objective is found within a tolerance of $1\times 10^{-6}$, and the number of iterations are averaged over 48 independent trials. Each image patch is 400 pixels in this figure.
In Fig. 6, the size of each image patch was chosen to be 400 pixels instead of 361 pixels for visualization purposes when the number of stored cost-to-goes is relatively small. When the network input is either a raw image patch or a complete representation of an image patch, a linear network will have exactly 400 inputs and 400 weight parameters, allowing it to store a maximum of 400 cost-to-go values. The curves for a $\times 1$ raw image and a $\times 1$ whitened image show these cases, and both curves get close to 400 stored cost-to-goes. However, input given by a whitened image generally takes far fewer least squares iterations to store the same number of cost-to-goes as a raw image, making it a computationally efficient image representation. The curve for a $\times 4$ raw image shows results for a natural image patch that has been upscaled by a factor of four using bicubic interpolation via Matlab’s imresize function. The curve for a $\times 4$ sparse code shows results for an overcomplete sparse code. Each of these representations has 1600 pixels (i.e., $4\times 400$ pixels, treating the coefficients of a Gabor function basis the same as double precision pixels in an image). Increasing the number of weights to 1600 allows a linear network to store a maximum of 1600 cost-to-go values. However, the best result for the $\times 4$ resized image remains below 400 stored cost-to-goes, and uses approximately the same number of least squares iterations as the best result for the $\times 4$ sparse code; which gets closer to 1600 stored cost-to-goes (the best result in Fig. 6 is 1323 cost-to-goes for 2701 least squares iterations).
These results confirm that a computationally efficient image representation must be both overcomplete and sparse in order to increase the storage capacity of a linear network. The key role of sparsity is discussed in loxley2, where it is shown how a sparse code decorrelates neighbouring pixels in an image and leads to an increase in the column-space dimension (rank) of the design matrix for the associated least squares problem. A sparse code also leads to a well-conditioned Hessian matrix, reducing the number of least squares iterations required for convergence. Since standard decorrelation techniques do not generalize to overcomplete representations, an overcomplete sparse code seems to be a minimum requirement for increasing neural network storage capacity efficiently.
Figure 7: Number of least squares iterations to store a number of cost-to-go values in a linear network using different image patch representations (see legend). The least squares objective is found within a tolerance of $1\times 10^{-6}$, and the number of iterations are averaged over 48 independent trials. Each image patch is 361 pixels.
Surprisingly, if we continue to increase the overcompleteness of image representations, the network storage capacity continues to increase. This is shown in Fig 7 for network inputs given by $\times 4$, $\times 16$, and $\times 64$ overcomplete sparse codes. The size of each image patch is 361 pixels; so the $\times 4$ sparse code has $4\times 361=1444$ pixels, the $\times 16$ sparse code has $16\times 361=5776$ pixels, and the $\times 64$ sparse code has $64\times 361=23104$ pixels. These quantities are shown as dashed vertical lines in Fig 7. The general result in Fig 7 is that the number of stored cost-to-goes (i.e., the network storage capacity) increases as the overcompleteness of an image representation increases. The first vertical dashed line at 361 cost-to-go values corresponds to the maximum storage capacity of a linear network with inputs given by complete codes. The other three vertical dashed lines correspond to maximum storage capacities for inputs given by $\times 4,\ \times 16,$ and $\times 64$ overcomplete sparse codes, respectively.
The best result for each representation in Fig 7 is summarized in Table 3. An extra data point is given in Table 3 to show the limit of our calculation at 21675 stored cost-to-go values (not shown in Fig 7). This limit holds because we have a maximum of 22201 image patches, while the parameterization of the benchmark set $D$ implemented here means the next data point is a benchmark with 22707 states – more than the available number of image patches.
Representation #\# LS Iterations #\# Stored Cost-To-Goes
$\times 1$ whitened image 177 243
$\times 4$ sparse code 2963 1323
$\times 16$ sparse code 4527 5043
$\times 64$ sparse code 9679 19683
$\times 64$ sparse code 20009 21675
Table 3: (Top) Four data points in Fig 7. (Bottom) One data point not shown in Fig  7.
Some variation in the average number of least-squares iterations can be seen in Table 3, particularly for the $\times 4$ sparse code. This is due to the relative proximity of each data point to its limiting case given by each of the vertical dashed lines. For example, the best result for the $\times 4$ sparse code is the data point at 1323 stored cost-to-go values, which is closer to its limiting value of 1444 than any other image representation is to its limiting value, and hence requires a relatively large number of least squares iterations as a result of the larger condition number of its Hessian matrix.
Now we are in a position to look at how the benchmark performs as the number of states increases. Increasing the number of states increases the size of the region over which a target may be tracked in the benchmark, as well as the number of stored cost-to-go values required for an optimal solution. Following from the results in Fig 7, we would expect that using efficient image representations should increase the number of states over which optimal target tracking is possible. Fig 8 shows this result: the expected total cost of tracking a target with $p=0.75$ over thirty time periods is shown for various image representations as the number of states of the benchmark is increased. The initial states are taken from Tables 1 and 2 and are $((0,1),(0,0))$ for the optimal tracker, and $((0,0),(0,0))$ for the greedy tracker. For comparison, exact results are shown for the expected total costs of optimal and greedy target tracking in the two right columns in Fig 8.
Figure 8: Expected total cost of tracking a target over 30 time periods with $p=0.75$ versus number of states of the benchmark, for different image-patch representations (see legend). Exact expected total costs for optimal and greedy target tracking (right columns) are shown for comparison.
According to Fig 8, a representation given by a $\times 1$ whitened image can solve the benchmark optimally for 243 states, but not for 363 states – as the expected total cost now exceeds that for greedy target tracking. The reason is, as in Fig 7, a complete code stores a maximum of 361 cost-to-go values which is less than 363. Similarly, a $\times 4$ sparse code can solve the benchmark optimally for 1323 states, but not for 1587 (>1444>1444) states; a $\times 16$ sparse code can solve the benchmark optimally for 5547 states, but not for 6075 (>5776>5776) states; and a $\times 64$ sparse code can solve the benchmark optimally for 21675 states. This was the largest number of image patches tested in this work. Actual values for the number of states shown in Fig 8 are due to the specific parameterization of the set $D$ implemented here.
4.5 Benchmark Performance for Different Initial States
We now look at the effect of starting at different initial states in the reinforcement learning benchmark. This is done by starting the benchmark in an initial state, and then running it with either an optimal or greedy policy until reaching the horizon. The total cost of each policy can then be compared. This is repeated for each state of the benchmark.
Training the benchmark with a $\times 64$ sparse code leads to 21675 possible initial states. The expected total cost of optimal and greedy policies is shown in Fig 9 for each initial state that leads to a suboptimal solution under the greedy policy. These initial states are indexed from 1 to 10880 and comprise roughly half of the 21675 available initial states. For the remaining 10795 initial states (not shown) the greedy policy is the optimal policy. The benchmark horizon is 200 time periods, and the Markov chain parameter is $p=0.4$. The policy found using fitted value iteration matches the optimal policy found using exact dynamic programming, as shown in Fig 9 (Inset) for initial states indexed from 4000 to 4050.
Figure 9: Expected total cost of optimal and greedy policies for each initial state leading to a suboptimal greedy policy. These initial states are indexed from 1 to 10880. “Fitted VI” is the policy found using fitted value iteration. Inset shows a close-up of initial states indexed from 4000 to 4050. The benchmark horizon is 200 time periods, and $p=0.4$ (see text for details).
Looking at a few specific initial states helps to understand Fig 9. The lowest cost optimal policy in Fig 9 starts at state $((0,0),(1,1))$ and has an expected total cost of 54. We can derive a theoretical bound for this state since it is part of the stationary policy in Table 1 (see also Figure 4). For $p=0$, its total cost is given by $(200-2)/3\times 1=66$. From Fig 4, we know the optimal cost for $p=0.4$ would be less than that for $p=0$, leading to $54<66$. We can also predict a value for the lowest cost greedy policy by making use of the infinite horizon result for the total cost ratio from the appendix: (1+1/(1p))(1+1/(1-p)). The expected total cost of the greedy policy can then be found by multiplying this ratio by the expected total cost of the optimal policy: for $p=0.4$ we get $(1+1/(1-p))\times 54=144$. This exactly matches the value in Fig 9, which is too small to see in that figure.
Now that we have some theoretical justification for the lowest cost policy in Fig 9, let us attempt to understand the highest cost policy in that figure. This policy starts at state ((42,42),(1,1))((42,-42),(1,1)) and has an expected total cost of $185870$ for the optimal policy, and $185890$ for the greedy policy. We can understand this as follows. Each number in the coordinate pair for $c_{k}-t_{k}\in D$ has the range [42,42][-42,42] when the number of states is 21675. Starting at state ((42,42),(1,1))((42,-42),(1,1)) means the tracker is as far behind the target as possible, but it is also as far to the right of the target as possible. Consider the corresponding situation when $p=0$: during three time periods, the target will move right once and up twice. If the tracker simply moves up three times over the same three time periods, then it has decreased the distance to the target by one every three time periods. Therefore, it can reach the target within $42\times 3=126$ time periods, and apply the stationary optimal policy thereafter. There is no other initial state with a suboptimal greedy policy that can have a larger expected total cost. The same argument holds approximately when $p=0.4$.
For the other 10795 initial states where the greedy policy is the optimal policy (not shown), the lowest cost policy starts at state $((0,0),(0,0))$ and has an expected total cost of 145. This is close to the expected total cost of the greedy policy for the initial state $((0,0),(1,1))$ just discussed. Predictably, the highest cost policy starts with the tracker as far behind the target as possible, in state ((42,42),(1,1))((-42,-42),(1,1)), and has an expected total cost of $701100$.
4.6 Benchmark Performance with Partial Training Sets
Now we consider situations where only some of the states (image patches) of the benchmark are used to train the network. One possibility would be to focus training on initial states that have a suboptimal greedy policy. However, under an optimal policy states outside this set will be visited. A better approach makes use of the analysis in the previous section to partition the benchmark states into two sets. States with at least one non-negative coordinate for $c_{k}-t_{k}$ in the benchmark correspond to a situation where the tracker is not both behind and to the left of the target. These states are independent of the set of states where the tracker is both behind and to the left of the target: in which case the tracker never reaches the target. When states partition in this way, it is possible to focus training on the relevant partition to achieve optimal performance.
Figure 10: Expected total cost of optimal and greedy policies for each initial state with a suboptimal greedy policy when only 78% of the training set is used (see Text for details). Initial states are indexed from 1 to 672. Fitted value iteration (Fitted VI) remains optimal (Inset). The benchmark horizon is 400 time periods, and $p=0.4$.
In Fig 10, the benchmark is used with a $\times 4$ sparse code leading to 1323 states (a small number of states is used here as it is easier to display). Of these, 1031 states have at least one non-negative coordinate (i.e., 78% of the total number of states). These states are used as the training set. Fitted value iteration is then applied to find optimal policies for the 672 initial states with suboptimal greedy policies. Fig 10 shows these policies remain optimal under the proposed training strategy. However, further decreasing the size of the training set rapidly degrades performance as states from the relevant partition start to be left out of the training set. The benchmark provides an ideal testing ground to investigate these types of behaviours.
5 Conclusion and Discussion
This work investigated optimal control over a sequence of natural images and provides three new contributions to reinforcement learning methodology.
The first consideration is what information to include in a state to allow for optimal control over a sequence of images. We found that an image can be a state provided certain conditions are met. The most obvious condition is that an image must contain all task-relevant information leading to a cost or reward at each time period. On the other hand, a less obvious condition is that any model of the dynamics must be capable of generating images from the environment. The reason is that a controllable Markov chain, or a Markov decision process, provides a mechanism for generating the next state from the current state and a chosen control. When each state is an image, the dynamics of moving through an environment generates a sequence of images.
The next consideration is how to encode an image, or find a suitable image representation, to make optimal control computationally efficient. Previous work suggested this might be possible by making use of sparse codes from visual neuroscience (loxley2). To test this idea we developed a new reinforcement learning benchmark that can be used to compare different image representations for computational efficiency during optimal control. The new benchmark is simple to implement, makes it easy to distinguish between optimal and suboptimal policies, and scales to large numbers of states and long horizons. The infinite horizon limit also leads to useful stationary policies. These properties are necessary to convincingly demonstrate the scaling of image representations and computational efficiency to large tasks with many states.
The main contribution of this work was to demonstrate that representing natural images as overcomplete sparse codes is computationally efficient for finding optimal policies in reinforcement learning. This is a significant result. The best result presented here was finding an optimal policy using a $\times 64$ overcomplete sparse code, allowing for the solution of an optimal control task approximately three orders of magnitude larger (in base 4) than could be solved using a complete code such as one generated using ICA or PCA. The underlying reason for the computational efficiency of an overcomplete sparse code can be related to the form of the linear least squares problem for training the final layer of a neural network. Provided the number of images is no larger than the size of an overcomplete sparse code, the design matrix of the least squares problem is approximately square. A key property of a sparse code is that neighbouring pixels in an input image are approximately decorrelated. Decorrelation increases the rank of the design matrix, and favourably conditions the Hessian matrix for neural network training (loxley2). In turn, this maximizes the storage capacity of the linear network and reduces the number of least squares iterations required for training. Since standard decorrelation techniques do not generalize to overcomplete representations, sparsity and overcompleteness seem to be minimum requirements for increasing neural network storage capacity efficiently. An interesting question then becomes: what is the limit to overcompleteness for a sparse code?
This work also shows deep learning is not necessary for efficient optimal control with natural images. A deep neural network may be able to learn specific features for a specific image dataset to help with optimal control. However, a sparse code only depends on the statistics of natural images, and therefore provides a set of general purpose features that can be used on any image dataset. Of course, deep learning can also be used to learn sparse codes (papyan; li). Our model has the advantage that it is simpler to train and uses less memory than a typical deep neural network. The time complexity of training our model depends upon the solution of two linear least squares problems. The first least squares problem has a $d\times m$ design matrix (for a single image with $d$ pixels and a sparse code of size $m$), leading to $O(dmt_{1})$ for an iterative algorithm such as gradient descent; with $t_{1}$ the number of iterations until convergence. For $n$ images this must be repeated $n$ times, leading to $O(dmt_{1}n)$. The second least squares problem has a $m\times n$ design matrix, leading to $O(mnt_{2})$; with $t_{2}$ the number of iterations until convergence. For optimal control over $N$ time periods this must be repeated $N$ times, leading to $O(mnt_{2}N)$. The overall time complexity of training is therefore $O(mn(dt_{1}+Nt_{2}))$. Both least squares problems are well-conditioned (at least until $n\rightarrow m$, see Figure 7) so the iteration numbers $t_{1}$ and $t_{2}$ are not excessive. Our model only requires memory to store $m(d+1)$ weights. Finally, it should be noted that sparse codes can also be generated directly from hardware using purpose built low-power neuromorphic computer chips (chip).
The original motivation for this work was to investigate possible computational advantages of the efficient coding hypothesis from neuroscience. If neural codes do in fact make use of sparseness and overcompleteness, this work has demonstrated some computational advantages to doing so. In particular, the decorrelating effect of a sparse code increases the speed of learning, while an overcomplete sparse code can additionally take advantage of an increase in the size and number of weights of a network to efficiently increase its memory capacity. These findings provide two possible explanations for the appearance of sparse codes in neuroscience.
Appendix A1: Method for generating overcomplete sparse codes
The method used here to generate overcomplete sparse codes of a natural images is taken from loxley1, loxley2, and starts with the (real-valued) two-dimensional (2D) Gabor function given by:
$G({r},{r}^{\prime})=A\exp{\left[-\frac{1}{2}\left(\frac{\tilde{i}^{2}}{\sigma_{x}({r}^{\prime})^{2}}+\frac{\tilde{j}^{2}}{\sigma_{y}({r}^{\prime})^{2}}\right)\right]}\cos{\left[k({r}^{\prime})\tilde{j}+\varphi({r}^{\prime})\right]},$ (19)
and
$(\tilde{i},\tilde{j})=\left(\begin{array}[]{cc}\cos{\phi({r}^{\prime})}&-\sin{\phi({r}^{\prime})}\\ \sin{\phi({r}^{\prime})}&\cos{\phi({r}^{\prime})}\\ \end{array}\right)\left(\begin{array}[]{c}i-x_{0}({r}^{\prime})\\ j-y_{0}({r}^{\prime})\\ \end{array}\right),$ (20)
where $k({r}^{\prime})=2\pi/\lambda({r}^{\prime})$; and where r=(i,j){r}=(i,j) and ${r}^{\prime}=(i^{\prime},j^{\prime})$ are discrete two-dimensional coordinates. When the 2D Gabor function is adapted to natural image statistics, the three spatial Gabor parameters $\sigma_{x},\sigma_{y},$ and $\lambda$ are found to be strongly correlated and have heavy-tailed distributions [loxley1]. The joint probability density of these parameter values is approximated using the sampling scheme in Table 4, and described by a Gaussian copula with Pareto marginal distributions.
Gabor Parameter(s) Sample Transformation
$\sigma_{x}^{\prime},\sigma_{y}^{\prime},\lambda^{\prime}$ $(\sigma_{x}^{\prime},\sigma_{y}^{\prime},\lambda^{\prime})=(1,1,\rho)z$
$\sigma_{x}$ $\sigma_{x}={\cal{PCDF}}^{-1}({\cal{NCDF}}(\sigma_{x}^{\prime}|0,1)|\alpha_{1},\beta_{1})$
$\sigma_{y}$ $\sigma_{y}={\cal{PCDF}}^{-1}({\cal{NCDF}}(\sigma_{y}^{\prime}|0,1)|\alpha_{2},\beta_{2})$
$\lambda$ $\lambda={\cal{PCDF}}^{-1}({\cal{NCDF}}(\lambda^{\prime}|0,1)|\alpha_{3},\beta_{3})$
Table 4: Sampling scheme for the three spatial Gabor function parameters: sample $z\sim{\cal{N}}(0,1)$ from the standard normal distribution, then apply the parameter transformations listed in the table. Here, $\rho,\alpha_{i}$, and $\beta_{i}$ are model parameters, ${\cal{PCDF}}^{-1}(x|\alpha,\beta)=\frac{\beta}{(1-x)^{1/\alpha}}$ is the inverse CDF for the Pareto distribution, and ${\cal{NCDF}}(x|0,1)$ denotes the CDF for the standard normal distribution.
Due to the Pareto marginal distributions the sampling scheme is length scale invariant. Scale invariance is a key property of natural images. The resulting set of randomly generated Gabor functions are therefore self-similar and multiscale in the same way as self-similar multiresolution wavelet schemes. All other Gabor parameters are sampled uniformly over their respective ranges.
In the first step, a sample is collected for each of the seven Gabor parameters $(\phi,\varphi,\sigma_{x},\sigma_{y},\lambda,x_{0},y_{0})$, leading to a single 2D Gabor function indexed by a value of ${r}^{\prime}$. This step is repeated $m$ times; leading to $m$ Gabor functions indexed by $m$ unique values of ${r}^{\prime}$. Two-dimensional Gabor functions are not orthogonal. However, given an image $I({r})\in\mathbb{R}^{d}$, it is possible to find its sparse code $a({r}^{\prime})\in\mathbb{R}^{m}$ using a least-squares approximation. Letting $G\in\mathbb{R}^{d\times m}$ be a matrix with elements $G({r},{r}^{\prime})$, the sparse code $a({r}^{\prime})$ is found by solving the least-squares problem,
$\text{minimize}\ \ \ \lVert Ga-I\rVert_{2}^{2}$
$\text{subject to}\ \ \ a\in\mathbb{R}^{m}.$
When $m>d$, the sparse code is overcomplete; meaning that there are more Gabor function coefficients than image pixels.
Appendix A2: Infinite horizon policy evaluation
A stationary policy is a policy that does not change with each time period $k$, and can therefore be written as $\pi=\{\mu(i),\mu(i),$...\}. In Sec. 4.1, we identified two stationary policies. We would now like to find the their expected total costs in the infinite horizon limit (i.e., as $N\rightarrow\infty$). To work in this limit, we must introduce a discount factor $0<\alpha\leq 1$ that prevents total costs from becoming infinite. For any stationary policy $\mu$, the expected total costs $J_{\mu}(1),...,J_{\mu}(n)$ are then found as the unique solution to the following system of linear equations:
$J_{\mu}(i)=g(i)+\alpha\sum_{j}p_{ij}(\mu(i))J_{\mu}(j).$ (21)
The stationary optimal policy identified in Sec. 4.1 is given by $\mu^{*}((0,1),(0,0))=(1,0)$, $\mu^{*}((0,0),(1,1))=(0,1)$, and $\mu^{*}((0,0),(0,1))=(0,1)$. These states have a cost per time period given by $g((0,1),(0,0))=1$, $g((0,0),(1,1))=0$, and $g((0,0),(0,1))=0$. After rearranging Eq. (21) as
$\sum_{j}\big(\delta_{ij}-\alpha p_{ij}(\mu(i))\big)J_{\mu}(j)=g(i),$ (22)
the linear system can be written as $\boldsymbol{AJ}_{\mu}=\boldsymbol{g}$, where
$\boldsymbol{A}=\begin{bmatrix}1&-\alpha&0\\ 0&1&-\alpha\\ -\alpha(1-p)&0&1-\alpha p\end{bmatrix}$, (23)
so that $\det{(\boldsymbol{A})}=1-\alpha p-\alpha^{3}(1-p)$. For the stationary optimal policy, $\boldsymbol{g}$ is given by $1,0,0]^{\top}$, and it is easily confirmed that $\boldsymbol{AJ}_{\mu}^{*}=[1,0,0]^{\top}$ is solved by
$\boldsymbol{J}_{\mu}^{*}=\begin{bmatrix}1-\alpha p\\ \alpha^{2}(1-p)\\ \alpha(1-p)\end{bmatrix}\times(1-\alpha p-\alpha^{3}(1-p))^{-1}.$ (24)
The stationary greedy policy identified in Sec. 4.1 is given by $\mu^{G}((0,0),(0,0))=(1,0)$, $\mu^{G}((0,-1),(1,1))=(0,1)$, and $\mu^{G}((0,-1),(0,1))=(0,1)$. These states have a cost per time period given by $g((0,0),(0,0))=0$, g((0,1),(1,1))=1g((0,-1),(1,1))=1, and g((0,1),(0,1))=1g((0,-1),(0,1))=1. Now $\boldsymbol{g}$ is given by $0,1,1]^{\top}$. It can then be confirmed that $\boldsymbol{AJ}_{\mu}^{G}=[0,1,1]^{\top}$ is solved by
$\boldsymbol{J}_{\mu}^{G}=\begin{bmatrix}\alpha+\alpha^{2}(1-p)\\ 1+\alpha(1-p)\\ 1+\alpha^{2}(1-p)\end{bmatrix}\times(1-\alpha p-\alpha^{3}(1-p))^{-1}.$ (25)
The expected total cost ratio $\boldsymbol{J}_{\mu}^{G}/\boldsymbol{J}_{\mu}^{*}$, for the first greedy state $i^{G}_{1}=(0,0),(0,0)$, and the first optimal state $i^{*}_{1}=(0,1),(0,0)$, is
$\frac{J_{\mu}^{G}(i^{G}_{1})}{J_{\mu}^{*}(i^{*}_{1})}=\alpha+\frac{\alpha^{2}}{1-\alpha p}.$ (26)
Taking the limit of zero discounting (i.e., $\alpha\rightarrow 1$), Eq. (26) becomes
$\lim_{\alpha\rightarrow 1}\frac{J_{\mu}^{G}(i^{G}_{1})}{J_{\mu}^{*}(i^{*}_{1})}=1+\frac{1}{1-p}.$ (27)
This ratio describes the gap in the expected total cost between the greedy and optimal policies in the infinite horizon limit, and is useful for comparing with the finite horizon results of Sec 4.1. Some useful values include: 2 (when $p=0$), 5 (when $p=0.75$), and 1/(1p)1/(1-p) as $p\rightarrow 1$. Alternatively, for $p=1$, the original expression (26) gives $1/(1-\alpha)$ as $\alpha\rightarrow 1$.
Appendix A3: Efficient evaluation of the expected total cost
An efficient method is now presented for evaluating the expected total cost,
${\mathbb{E}}\left\{\sum_{k=0}^{N-1}g_{k}(i_{k},\mu_{k}(i_{k}),i_{k+1})\right\}.$
The key to an efficient evaluation is to make use of the Markov property to decompose the joint probability distribution $p(i_{0},i_{1},...,i_{N})$ into a product of transition probabilities: $p_{i_{0}}p_{i_{0}i_{1}}p_{i_{1}i_{2}}...p_{i_{N-1}i_{N}}$; where each transition probability $p_{i_{k}i_{k+1}}(\mu_{k}(i_{k}))$ has been abbreviated as $p_{i_{k}i_{k+1}}$. It is then possible to write the expected total cost as:
${\mathbb{E}}\left\{\sum_{k=0}^{N-1}g_{k}(i_{k},\mu_{k}(i_{k}),i_{k+1})\right\}$
$=\sum_{i_{0}}p_{i_{0}}\sum_{i_{1}}p_{i_{0}i_{1}}g_{0}(i_{0},\mu_{0}(i_{0}),i_{1})$
$\ \ \ \ +\sum_{i_{0}}p_{i_{0}}\sum_{i_{1}}p_{i_{0}i_{1}}\sum_{i_{2}}p_{i_{1}i_{2}}g_{1}(i_{1},\mu_{1}(i_{1}),i_{2})$
$\ \ \ \ +...+\sum_{i_{0}}p_{i_{0}}...\sum_{i_{N-1}}p_{i_{N-2}i_{N-1}}\sum_{i_{N}}p_{i_{N-1}i_{N}}g_{N-1}(i_{N-1},\mu_{N-1}(i_{N-1}),i_{N}).$
The number of evaluations here is formidable: the last term is a product of $N$ sums, each over $|S|$ states; leading to $|S|^{N}$ evaluations. We can do much better than this by taking advantage of the Markov property of the transition probabilities. Defining the following recurrence relation:
$f_{i_{0}}$ $=p_{i_{0}},$ (28)
$f_{i_{k}}$ $=\sum_{i_{k-1}}f_{i_{k-1}}p_{i_{k-1}i_{k}},$ (29)
allows the previous equation to be written as:
${\mathbb{E}}\left\{\sum_{k=0}^{N-1}g_{k}(i_{k},\mu_{k}(i_{k}),i_{k+1})\right\}$
$=\sum_{i_{0}}f_{i_{0}}\sum_{i_{1}}p_{i_{0}i_{1}}g_{0}(i_{0},\mu_{0}(i_{0}),i_{1})$
$\ \ \ \ +\sum_{i_{1}}f_{i_{1}}\sum_{i_{2}}p_{i_{1}i_{2}}g_{1}(i_{1},\mu_{1}(i_{1}),i_{2})$
$\ \ \ \ +...+\sum_{i_{N-1}}f_{i_{N-1}}\sum_{i_{N}}p_{i_{N-1}i_{N}}g_{N-1}(i_{N-1},\mu_{N-1}(i_{N-1}),i_{N}),$
$=\sum_{k=0}^{N-1}\sum_{i_{k}}f_{i_{k}}\sum_{i_{k+1}}p_{i_{k}i_{k+1}}g_{k}(i_{k},\mu_{k}(i_{k}),i_{k+1}).$
Now we only need to perform $O(N|S|^{2})$ evaluations given by the product of two sums over $|S|$ states, followed by a sum over $N$ time periods. However, when the number of states $|S|$ is large, $O(N|S|^{2})$ evaluations is still inefficient. A further simplification of the total expected cost results for our particular form of the cost per time period $g_{k}$, leading to:
${\mathbb{E}}\left\{\sum_{k=0}^{N-1}g_{k}(i_{k})\right\}=\sum_{k=0}^{N-1}\sum_{i_{k}}f_{i_{k}}g_{k}(i_{k}).$ (30)
Upon defining $i_{k-1}=(a_{1},b_{1})$ and $i_{k}=(a_{2},b_{2})$, Eq. (29) can be written as
$f_{(a_{2},b_{2})}=\sum_{a_{1},b_{1}}f_{(a_{1},b_{1})}p(b_{2}|b_{1})\delta(a_{2}-(a_{1}+\mu(a_{1},b_{1})-b_{2})),$
where Eq. (15) was used for $p_{i_{k-1}i_{k}}$. Multiplying by $g_{k}(i_{k})$, and summing over $i_{k}=(a_{2},b_{2})$, then leads to
$\sum_{i_{k}}f_{i_{k}}g_{k}(i_{k})=\sum_{a_{1},b_{1},b_{2}}f_{(a_{1},b_{1})}p(b_{2}|b_{1})g_{k}(a_{1}+\mu(a_{1},b_{1})-b_{2},b_{2}).$
Since $a_{1}\in D$ and $b_{1},b_{2}\in T$, this sum is $O(|S|)$ because the number of states can be written as $|S|=|D||T|$; and $|T|=3$. Using this sum in Eq. (30), we see that the total number of evaluations required to determine the expected total cost is $O(N|S|)$.